Consider light with a wavelength of 519 nm is passing through a diffraction grating with 1208 lines per cm. At what distance from the central bright maximum will the 1st order maximum be seen on a screen placed 1.20 m from the diffraction grating. Give your answer in centimeters (cm) with 3 significant figures please. :)
Answer #1

\lambda = wavelength = 519 nm = 519 x 10-9 m

d = slit width = 0.01/1208 = 8.3 x 10-6 m

D = screen distance = 1.20 m

n = 1

y = distance from central bright maximum

using the equation

y = nD\lambda/d

y = 1 (1.20) (519 x 10-9) /(8.3 x 10-6) = 0.075 m = 7.50 cm